Oranges chosen at random have a mean mass of 210 grams and standard deviation of 35 grams. Assuming a normal distribution, what is the probability of choosing an orange with a mass between 170 grams and 250 grams

1
Expert's answer
2022-01-25T17:33:33-0500

Given: "\mu=210,\sigma=35"

"P(170 < X < 250)=P(X<250)-P(X<170)"

Now,

"\begin{aligned}
& P[x <250] \\
=& P\left[z <\frac{X-\mu}{\sigma}\right] \\
=& P\left[z <\frac{250-210}{35}\right] \\
=& P[z <1.1428] \\
=& 0.87344 \\
\end{aligned}"

and

"\begin{aligned}
& P[x <170] \\
=& P\left[z <\frac{X-\mu}{\sigma}\right] \\
=& P\left[z <\frac{170-210}{35}\right] \\
=& P[z <-1.1428] \\
=& 0.12656 \\
\end{aligned}"

So,

"P(170<X<250)=P(X<250)\u2212P(X<170)=0.87344-0.12656=0.74688"